OpenStudy (anonymous):

Ok, I officially suck at this. Help? What is the equation of the line in slope-intercept form that passes through the point (3, –5) and is perpendicular to the line y = 3x – 6? y = –3x – 4 y = –3x + 4 y = -1/3x – 4 y = -1/3x + 4

OpenStudy (anonymous):

You need to apply the point slope equation for this to be able to solve it but you also need to remember a little thing about how you can know the slope of a line perpendicular to another line that you know the slope of

OpenStudy (anonymous):

Yeah, Idk how to do that.

OpenStudy (anonymous):

Do you remember what the slope is of a line perpendicular to another line that has a known slope?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

its the opposite inverse of the slope of the line it is perpendicular to

OpenStudy (anonymous):

first take your (3,-5), and put -5, into y and 3 into x

OpenStudy (anonymous):

-5=3(3)-6

OpenStudy (anonymous):

> opposite inverse Negative reciprocal is clearler

OpenStudy (anonymous):

You guys are confusing me.

OpenStudy (anonymous):

A line perpendicular to another line has a slope that is the negative reciprocal of the original slope. For example: y = 4x + 7 --> The slope of this line is m = 4. But the slope of the line perpendicular to this one would be the negative reciprocal. A reciprocal is just a number or fraction flipped up-side down. Like 4 becomes 1/4 or -2/3 becomes 3/-2 and so on. In this case, the perpendicular line has a negative reciprocal slope of 4, which is -1/4. This is because the reciprocal of 4 is 1/4. When you make that negative, it becomes - 1/4. Pretty easy right? So now, the question here asks us to give the equation of a line perpendicular to y = 3x - 6 and goes through (3, -5). First we know the slope of this perpendicular line will be the negative reciprocal of 3 which is - 1/3. So the equation will look something like y = -1/3x + b, we don't know what b is yet. However, since we have a point (3, -5) given and the question tells us that the perpendicular line goes through this point, then we can substitute that point in to our y = -1/3x + b equation and solve for b. In the point (3, -5), 3 is the x-value, and -5 is the y-value. Substitute these in and solve for b. -5 = -1/3(3) + b -5 = -1 + b --> The 3's cancel out when we multiply the fraction -1/3 by 3. -5 + 1 = b --> Bring -1 over to other side. -4 = b --> Now we know what b so we can complete our slope-intercept equation of this line. Therefore, the equation in slope-intercept form is: y = -1/3x - 4 Where -4 = b and -1/3 is the slope. @britbrat1717

OpenStudy (anonymous):

@genius12 Thank you!